OpenStudy (anonymous):

What is the sum of the geometric sequence 8, -16, 32 . if there are 15 terms?

OpenStudy (anonymous):

ooh careful on this one now it is geometric not aritmatic and it is a sum not a sequence :P

OpenStudy (anonymous):

yesss ic this lol but these are the ones that im just like....... kill me now lol

OpenStudy (anonymous):

ar0+ar1+ar2 so a is 8 now we try to find r by looking at its 2nd term -16=8*r sfr r=-16/8 which is -2 right? with me so far?

OpenStudy (anonymous):

yup :P

OpenStudy (anonymous):

to find sum we use a*r^(n+1)/(r-1)-a/(r-1) that is just a formula n is how many terms we want to add up but we start at 0 you want 15 terms so we subtract 1 so n=15 a=8 and r=-2 :) just plug in the numbers

OpenStudy (anonymous):

87384